# Measurement of the power spectrum turnover scale from the cross-correlation between CMB lensing and Quaia

### v2

@@ 4.2 Measurement of the turnover scale
Table 1: Constraints on the logarithm of the stretch parameter $\alpha$ characterising the position of the power spectrum turnover for different analysis choices. Details are discussed in the main text.
@@
The stretch parameter $\alpha$ quantifies a potential shift in the scale dependence of the matter power spectrum with respect to our fiducial model. Although the most prominent feature that can in principle drive the constraints on $\alpha$ is the turnover, both the broadband curvature of the power spectrum on scales smaller than this turnover and the BAO wiggles can also act as anchors and effective standard rulers. It is worth noting that this power spectrum curvature has been indeed used in recent works as an effective standard ruler to constrain cosmological parameters, particularly $H_{0}$ (Philcox et al., 2021; Farren et al., 2022; Philcox et al., 2022). Treating the small-scale slope $\gamma$ as a free parameter should minimise our dependence on the power spectrum curvature, and indeed we see in Fig. 9 that there is a clear degeneracy between $\gamma$ and $\alpha$. Thus it is worth quantifying how much constraining power we lose by discarding this information. Repeating our analysis fixing $\gamma$ to its fiducial value $\gamma=1$ (row 16) results in constraints on $\log_{10}\alpha$ that are indeed significantly tighter: $\log_{10}\alpha=-0.082^{+0.051}_{-0.049}$, corresponding to a factor $\sim 2$ smaller uncertainties, and a difference of $\sim 1.6\sigma$ with respect to the fiducial position of the turnover scale ($\alpha=1$).
@@ 5 Conclusions
These measurements translate into a $\sim 20\%$ determination of the comoving scale of the turnover. The measurement is completely driven by the turnover itself, and not from the curvature of the power spectrum on smaller scales, used in previous studies to obtain improved cosmological constraints. Assuming that the small-scale dependence of the power spectrum is known leads to a factor of $\sim 2$ reduction in the error on the turnover scale. Assuming the cosmological parameters preferred by Planck data, the inferred equality scale $k_{\rm eq}=0.0093^{+0.0025}_{-0.0021}\,{\rm Mpc}^{-1}$ is in reasonable agreement with the value measured from CMB data. The measurement of the turnover scale is highly robust to the potential presence of large- and small-scale systematics, the mis-modelling of quasar bias, and to various analysis choices (see Table 1). Furthermore, we have shown that, as a standard ruler, our measurements are completely driven by the power spectrum turnover, and receive no significant information from the BAO wiggles.
@@ Acknowledgements
We thank Pedro Ferreira, Carlos García-García, Maria Vincenzi, and Sebastian von-Hausseger for useful comments and discussions. DA acknowledges support from UKSA under grant ST/Y005902/1 and from the Beecroft Trust. GF acknowledges the support of the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 851274). We made extensive use of computational resources at the University of Oxford Department of Physics, funded by the John Fell Oxford University Press Research Fund.

### v3

@@ 4.2 Measurement of the turnover scale
$17.\,{\rm Fixed}\,\,\beta$ $-0.069^{+0.081}_{-0.056}$
Table 1: Constraints on the logarithm of the stretch parameter $\alpha$ characterising the position of the power spectrum turnover for different analysis choices. These results are also presented in Fig. 10, and further details are discussed in the main text.
@@
The stretch parameter $\alpha$ quantifies a potential shift in the scale dependence of the matter power spectrum with respect to our fiducial model. Although the most prominent feature that can in principle drive the constraints on $\alpha$ is the turnover, both the broadband curvature of the power spectrum on scales smaller than this turnover and the BAO wiggles can also act as anchors and effective standard rulers. It is worth noting that this power spectrum curvature has been indeed used in recent works as an effective standard ruler to constrain cosmological parameters, particularly $H_{0}$ (Philcox et al., 2021; Farren et al., 2022; Philcox et al., 2022). Treating the small-scale slope $\gamma$ as a free parameter should minimise our dependence on the power spectrum curvature, and indeed we see in Fig. 9 that there is a clear degeneracy between $\gamma$ and $\alpha$. Thus it is worth quantifying how much constraining power we lose by suppressing this information. Repeating our analysis fixing $\gamma$ to its fiducial value $\gamma=1$ (row 16) results in constraints on $\log_{10}\alpha$ that are indeed significantly tighter: $\log_{10}\alpha=-0.082^{+0.051}_{-0.049}$, corresponding to a factor $\sim 2$ smaller uncertainties, and a difference of $\sim 1.6\sigma$ with respect to the fiducial position of the turnover scale ($\alpha=1$). In contrast, fixing the large-scale slope to $\beta=1$ instead (see row 17 in Table 1) leads to a milder $\sim 25\%$ decrease in the error on $\log_{10}\alpha$, and a less-significant $\sim 1\sigma$ downwards shift in its best-fit value (see Fig. 10). Thus, although both assumptions about the power spectrum slope on scales above and below the turnover play a significant role in determining its position, the impact of the better-measured small-scale slope is larger. The recovered value of $\gamma$ in our fiducial analysis is $\gamma=0.82\pm 0.16$, compatible with 1 within $1.1\sigma$. Fixing $\beta=1$ does not change the uncertainty with which $\gamma$ is determined, since constraints on both parameters are driven by different scale ranges, but it pushes its preferred value $\sim 0.3\sigma$ closer to $1$: $\gamma=0.87\pm 0.16$. Our preferred model is therefore compatible with the $\Lambda$CDM expectation.
@@ 5 Conclusions
These measurements translate into a $\sim 20\%$ determination of the comoving scale of the turnover. The measurement is driven by the turnover itself, and not from the curvature of the power spectrum on smaller scales, used in previous studies to obtain improved cosmological constraints. Assuming that the small-scale dependence of the power spectrum is known leads to a factor of $\sim 2$ reduction in the error on the turnover scale. Assuming the cosmological parameters preferred by Planck data, the inferred equality scale $k_{\rm eq}=0.0093^{+0.0025}_{-0.0021}\,{\rm Mpc}^{-1}$ is in reasonable agreement with the value measured from CMB data. The measurement of the turnover scale is highly robust to the potential presence of large- and small-scale systematics, the mis-modelling of quasar bias, and to various analysis choices (see Table 1). Furthermore, we have shown that, as a standard ruler, our measurements are completely driven by the power spectrum turnover, and receive no significant information from the BAO wiggles.
@@ Acknowledgements
We thank Pedro Ferreira, Carlos García-García, Maria Vincenzi, Sebastian von-Hausseger, and the anonymous referee for useful comments and discussions. DA acknowledges support from UKSA under grant ST/Y005902/1 and from the Beecroft Trust. GF acknowledges the support of the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 851274). We made extensive use of computational resources at the University of Oxford Department of Physics, funded by the John Fell Oxford University Press Research Fund.